# A Practical Examination of AI-Generated Text Detectors for Large Language Models

### v1

@@ 2 Related Work and Background
Evaluated Detectors and Datasets. In our paper, we evaluate six representative detectors: RADAR (Hu et al., 2023), Detection in the Wild (Wild) (Li et al., 2024), T5Sentinel (Chen et al., 2023), Fast-DetectGPT (Bao et al., 2024), GPTID (Tulchinskii et al., 2023), LogRank (Mitchell et al., 2023), and Binoculars (Hans et al., 2024). RADAR, Wild, and T5Sentinel are trained detectors, while Fast-DetectGPT, GPTID, LogRank, and Binoculars are zero-shot detectors. To ensure a fair comparison and assess the detectors’ ability to generalize to new data, we carefully select datasets that have not been used in the training or evaluation of these detectors. Table 1 presents an overview of the datasets and domains on which each detector has been evaluated. Several datasets, such as Xsum, SQuAD, and Reddit Writing Prompts, have been used in the evaluation or training of multiple detectors. Although these detectors achieve strong Area Under the Receiver Operating Characteristic (AUROC) scores on these datasets, they do not report the True Positive Rate at a set False Positive Rate (TPR@FPR), which is a crucial metric in real-world scenarios. To address this gap, we aim to evaluate all six detectors on the same datasets using both AUROC and TPR at FPR metrics.
@@ 6 Conclusion
This study evaluates six advanced detectors across seven tasks and four languages, revealing notable inconsistencies in their detection capabilities. We also examined three different prompting strategies and their impact on detectability, finding that requests for more "human-like" output do not make the text harder to detect, while rewritten human content proves more difficult to identify.
@@ A.1 Results by Model
| GPT-4o | Radar | 0.00 | 0.6829 |
| Fast-DetectGPT | 0.05 | 0.9013 | |
| Wild | 0.00 | 0.3936 | |
| PHD | 0.00 | 0.6308 | |
| LogRank | 0.00 | 0.4573 | |
| T5Sentinel | 0.00 | 0.4288 | |
| Binoculars | 0.00 | 0.0721 | |
| Llama-3 | Radar | 0.03 | 0.8287 |
| Fast-DetectGPT | 0.26 | 0.8602 | |
| Wild | 0.03 | 0.5504 | |
| PHD | 0.00 | 0.5822 | |
| LogRank | 0.00 | 0.4646 | |
| T5Sentinel | 0.00 | 0.4844 | |
| Binoculars | 0.00 | 0.0102 | |
| Mistral | Radar | 0.00 | 0.5444 |
| Fast-DetectGPT | 0.15 | 0.9127 | |
| Wild | 0.00 | 0.4815 | |
| PHD | 0.00 | 0.7464 | |
| LogRank | 0.00 | 0.5925 | |
| T5Sentinel | 0.02 | 0.6851 | |
| Binoculars | 0.00 | 0.0157 | |
| Phi-3 | Radar | 0.01 | 0.7630 |
| Fast-DetectGPT | 0.29 | 0.9512 | |
| Wild | 0.00 | 0.4487 | |
| PHD | 0.00 | 0.6597 | |
| LogRank | 0.00 | 0.5176 | |
| T5Sentinel | 0.00 | 0.4498 | |
| Binoculars | 0.00 | 0.0581 |
@@
| GPT-4o | Radar | 0.03 | 0.3139 |
| Fast-DetectGPT | 0.47 | 0.9468 | |
| Wild | 0.03 | 0.5078 | |
| PHD | 0.02 | 0.5238 | |
| LogRank | 0.00 | 0.4098 | |
| T5Sentinel | 0.04 | 0.4996 | |
| Binoculars | 0.00 | 0.0684 | |
| Llama-3 | Radar | 0.15 | 0.6888 |
| Fast-DetectGPT | 0.85 | 0.9865 | |
| Wild | 0.08 | 0.6404 | |
| PHD | 0.00 | 0.2484 | |
| LogRank | 0.00 | 0.1682 | |
| T5Sentinel | 0.01 | 0.5277 | |
| Binoculars | 0.00 | 0.0199 | |
| Mistral | Radar | 0.06 | 0.6091 |
| Fast-DetectGPT | 0.77 | 0.9626 | |
| Wild | 0.01 | 0.5764 | |
| PHD | 0.00 | 0.3239 | |
| LogRank | 0.00 | 0.2548 | |
| T5Sentinel | 0.01 | 0.4764 | |
| Binoculars | 0.00 | 0.0330 | |
| Phi-3 | Radar | 0.07 | 0.6323 |
| Fast-DetectGPT | 0.58 | 0.9183 | |
| Wild | 0.12 | 0.6421 | |
| PHD | 0.00 | 0.2517 | |
| LogRank | 0.00 | 0.1543 | |
| T5Sentinel | 0.01 | 0.3761 | |
| Binoculars | 0.00 | 0.1322 |
@@
| GPT-4o | Radar | 0.00 | 0.1771 |
| Fast-DetectGPT | 0.14 | 0.7731 | |
| Wild | 0.15 | 0.5088 | |
| PHD | 0.00 | 0.4938 | |
| LogRank | 0.00 | 0.2668 | |
| T5Sentinel | 0.03 | 0.5675 | |
| Binoculars | 0.00 | 0.2634 | |
| Llama-3 | Radar | 0.01 | 0.5740 |
| Fast-DetectGPT | 0.23 | 0.7756 | |
| Wild | 0.20 | 0.7097 | |
| PHD | 0.00 | 0.2203 | |
| LogRank | 0.00 | 0.0879 | |
| T5Sentinel | 0.04 | 0.5902 | |
| Binoculars | 0.00 | 0.1417 | |
| Mistral | Radar | 0.00 | 0.3100 |
| Fast-DetectGPT | 0.10 | 0.5555 | |
| Wild | 0.20 | 0.6626 | |
| PHD | 0.00 | 0.4494 | |
| LogRank | 0.00 | 0.2037 | |
| T5Sentinel | 0.07 | 0.5729 | |
| Binoculars | 0.00 | 0.4232 | |
| Phi-3 | Radar | 0.16 | 0.7253 |
| Fast-DetectGPT | 0.19 | 0.5581 | |
| Wild | 0.53 | 0.8998 | |
| PHD | 0.00 | 0.0829 | |
| LogRank | 0.05 | 0.1406 | |
| T5Sentinel | 0.03 | 0.5877 | |
| Binoculars | 0.36 | 0.5467 |
@@
| GPT-4o | Radar | 0.05 | 0.6134 |
| Fast-DetectGPT | 0.69 | 0.9712 | |
| Wild | 0.00 | 0.5466 | |
| PHD | 0.00 | 0.3374 | |
| LogRank | 0.00 | 0.2094 | |
| T5Sentinel | 0.06 | 0.4982 | |
| Binoculars | 0.00 | 0.0037 | |
| Llama-3 | Radar | 0.14 | 0.6941 |
| Fast-DetectGPT | 0.82 | 0.9824 | |
| Wild | 0.00 | 0.6489 | |
| PHD | 0.00 | 0.2356 | |
| LogRank | 0.00 | 0.1575 | |
| T5Sentinel | 0.01 | 0.4588 | |
| Binoculars | 0.00 | 0.0019 | |
| Mistral | Radar | 0.31 | 0.8296 |
| Fast-DetectGPT | 0.64 | 0.9532 | |
| Wild | 0.00 | 0.6000 | |
| PHD | 0.01 | 0.2621 | |
| LogRank | 0.00 | 0.2279 | |
| T5Sentinel | 0.01 | 0.4856 | |
| Binoculars | 0.00 | 0.0157 | |
| Phi-3 | Radar | 0.06 | 0.7621 |
| Fast-DetectGPT | 0.75 | 0.9066 | |
| Wild | 0.22 | 0.7177 | |
| PHD | 0.00 | 0.1411 | |
| LogRank | 0.00 | 0.0410 | |
| T5Sentinel | 0.00 | 0.2760 | |
| Binoculars | 0.00 | 0.0045 |
@@
| GPT-4o | Radar | 0.00 | 0.2464 |
| Fast-DetectGPT | 0.46 | 0.9547 | |
| Wild | 0.03 | 0.6328 | |
| PHD | 0.03 | 0.7828 | |
| LogRank | 0.00 | 0.3659 | |
| T5Sentinel | 0.02 | 0.6542 | |
| Binoculars | 0.00 | 0.0540 | |
| Llama-3 | Radar | 0.41 | 0.8593 |
| Fast-DetectGPT | 0.91 | 0.9847 | |
| Wild | 0.57 | 0.9379 | |
| PHD | 0.00 | 0.3000 | |
| LogRank | 0.00 | 0.0858 | |
| T5Sentinel | 0.00 | 0.1860 | |
| Binoculars | 0.00 | 0.0037 | |
| Mistral | Radar | 0.00 | 0.1914 |
| Fast-DetectGPT | 0.48 | 0.9385 | |
| Wild | 0.04 | 0.5701 | |
| PHD | 0.00 | 0.6529 | |
| LogRank | 0.00 | 0.2980 | |
| T5Sentinel | 0.00 | 0.4285 | |
| Binoculars | 0.00 | 0.0835 | |
| Phi-3 | Radar | 0.55 | 0.9030 |
| Fast-DetectGPT | 0.55 | 0.7656 | |
| Wild | 0.41 | 0.9322 | |
| PHD | 0.00 | 0.3058 | |
| LogRank | 0.02 | 0.1644 | |
| T5Sentinel | 0.08 | 0.6950 | |
| Binoculars | 0.00 | 0.3388 |
@@
| GPT-4o | Radar | 0.14 | 0.9800 |
| Fast-DetectGPT | 0.98 | 0.9986 | |
| Wild | 0.00 | 0.9844 | |
| PHD | 0.00 | 0.1179 | |
| LogRank | 0.00 | 0.0109 | |
| T5Sentinel | 0.00 | 0.2526 | |
| Binoculars | 0.00 | 0.0000 | |
| Llama-3 | Radar | 0.53 | 0.9701 |
| Fast-DetectGPT | 0.97 | 0.9870 | |
| Wild | 0.44 | 0.9933 | |
| PHD | 0.00 | 0.0221 | |
| LogRank | 0.00 | 0.0061 | |
| T5Sentinel | 0.00 | 0.3243 | |
| Binoculars | 0.00 | 0.0090 | |
| Mistral | Radar | 0.44 | 0.9830 |
| Fast-DetectGPT | 1.00 | 0.9990 | |
| Wild | 0.55 | 0.9948 | |
| PHD | 0.00 | 0.0734 | |
| LogRank | 0.00 | 0.0080 | |
| T5Sentinel | 0.00 | 0.2331 | |
| Binoculars | 0.00 | 0.0000 | |
| Phi-3 | Radar | 0.64 | 0.9036 |
| Fast-DetectGPT | 0.70 | 0.7981 | |
| Wild | 0.44 | 0.9818 | |
| PHD | 0.00 | 0.0574 | |
| LogRank | 0.06 | 0.1677 | |
| T5Sentinel | 0.01 | 0.4661 | |
| Binoculars | 0.03 | 0.1535 |
@@
| GPT-4o | Radar | 0.02 | 0.6151 |
| Fast-DetectGPT | 0.05 | 0.6543 | |
| Wild | 0.11 | 0.5532 | |
| PHD | 0.00 | 0.4309 | |
| LogRank | 0.00 | 0.4135 | |
| T5Sentinel | 0.03 | 0.5792 | |
| Binoculars | 0.00 | 0.3095 | |
| Llama-3 | Radar | 0.13 | 0.9048 |
| Fast-DetectGPT | 0.39 | 0.7758 | |
| Wild | 0.37 | 0.7109 | |
| PHD | 0.00 | 0.2959 | |
| LogRank | 0.00 | 0.2991 | |
| T5Sentinel | 0.02 | 0.4761 | |
| Binoculars | 0.00 | 0.1452 | |
| Mistral | Radar | 0.02 | 0.8374 |
| Fast-DetectGPT | 0.09 | 0.6568 | |
| Wild | 0.19 | 0.6400 | |
| PHD | 0.00 | 0.3708 | |
| LogRank | 0.01 | 0.4333 | |
| T5Sentinel | 0.03 | 0.5486 | |
| Binoculars | 0.01 | 0.2780 | |
| Phi-3 | Radar | 0.18 | 0.9735 |
| Fast-DetectGPT | 0.06 | 0.5191 | |
| Wild | 0.44 | 0.7872 | |
| PHD | 0.00 | 0.2557 | |
| LogRank | 0.02 | 0.4431 | |
| T5Sentinel | 0.09 | 0.5473 | |
| Binoculars | 0.00 | 0.3240 |
@@ A.3 Results by Detector
| Code | Reviews | Abstracts | Translation ES | Translation FR | Translation ZH | |||||||
@@
| Radar | 0.0100 | 0.7047 | 0.4375 | 0.9592 | 0.2475 | 0.5500 | 0.0000 | 0.8179 | 0.0875 | 0.8372 | 0.4000 | 0.8282 |
| Fast-DetectGPT | 0.1875 | 0.9064 | 0.9125 | 0.9457 | 0.6000 | 0.9109 | 0.0000 | 0.5569 | 0.0150 | 0.5406 | 0.3275 | 0.7519 |
| Wild | 0.0075 | 0.4686 | 0.3575 | 0.9886 | 0.2800 | 0.7682 | 0.0000 | 0.6981 | 0.0175 | 0.5025 | 0.6275 | 0.9093 |
| PHD | 0.0075 | 0.3452 | 0.2450 | 0.9323 | 0.1125 | 0.4896 | 0.0000 | 0.5203 | 0.0050 | 0.6322 | 0.3325 | 0.8556 |
| LogRank | 0.0025 | 0.4920 | 0.6600 | 0.9518 | 0.0775 | 0.7715 | 0.0000 | 0.3291 | 0.0275 | 0.5157 | 0.1825 | 0.7833 |
| T5Sentinel | 0.0425 | 0.4880 | 0.0850 | 0.6810 | 0.0800 | 0.5091 | 0.0000 | 0.2807 | 0.0000 | 0.3361 | 0.0900 | 0.6437 |
| Binoculars | 0.4175 | 0.9610 | 0.9325 | 0.9594 | 0.6625 | 0.8800 | 0.0000 | 0.6382 | 0.0525 | 0.6104 | 0.4850 | 0.8529 |
@@
| Radar | 0.1725 | 0.8101 | 0.0125 | 0.4907 | 0.0600 | 0.5478 | 0.0100 | 0.3266 |
| Fast-DetectGPT | 0.7300 | 0.9564 | 0.5525 | 0.9419 | 0.8325 | 0.9776 | 0.7400 | 0.9543 |
| Wild | 0.1625 | 0.7964 | 0.0575 | 0.5149 | 0.0250 | 0.4940 | 0.0325 | 0.6434 |
| PHD | 0.0400 | 0.4496 | 0.1500 | 0.7072 | 0.0175 | 0.7027 | 0.1175 | 0.8104 |
| LogRank | 0.0475 | 0.7835 | 0.0000 | 0.7088 | 0.0200 | 0.8103 | 0.2000 | 0.8775 |
| T5Sentinel | 0.0600 | 0.6184 | 0.0000 | 0.4281 | 0.0000 | 0.4337 | 0.0075 | 0.7062 |
| Binoculars | 0.9300 | 0.9764 | 0.7650 | 0.8903 | 0.9450 | 0.9860 | 0.6500 | 0.9291 |
@@
| Radar | 0.0875 | 0.5822 | 0.0200 | 0.3396 | 0.0275 | 0.3111 | 0.1825 | 0.6689 |
| Fast-DetectGPT | 0.2525 | 0.8434 | 0.0950 | 0.6061 | 0.0875 | 0.5549 | 0.1450 | 0.6529 |
| Wild | 0.1525 | 0.7993 | 0.2800 | 0.6319 | 0.2475 | 0.5853 | 0.5975 | 0.9791 |
| PHD | 0.1900 | 0.6191 | 0.2675 | 0.6989 | 0.2950 | 0.6680 | 0.3500 | 0.7762 |
| LogRank | 0.7800 | 0.9594 | 0.4200 | 0.9019 | 0.4050 | 0.8821 | 0.5750 | 0.9114 |
| T5Sentinel | 0.1200 | 0.6325 | 0.0000 | 0.2780 | 0.0000 | 0.3037 | 0.0125 | 0.3730 |
| Binoculars | 0.4050 | 0.8027 | 0.1875 | 0.6270 | 0.1250 | 0.5827 | 0.0600 | 0.6127 |
@@
| Radar | 0.8700 | 0.9955 | 0.1000 | 0.7255 | 0.0675 | 0.6067 | 0.0600 | 0.5850 |
| Fast-DetectGPT | 0.8925 | 0.9902 | 0.5350 | 0.9368 | 0.7700 | 0.9500 | 0.5300 | 0.9258 |
| Wild | 0.0300 | 0.8969 | 0.1075 | 0.6607 | 0.1475 | 0.6448 | 0.0975 | 0.2812 |
| PHD | 0.0725 | 0.6497 | 0.2725 | 0.8254 | 0.2700 | 0.8235 | 0.5945 | 0.9209 |
| LogRank | 0.9625 | 0.9983 | 0.4600 | 0.9055 | 0.4425 | 0.9056 | 0.6125 | 0.8754 |
| T5Sentinel | 0.1775 | 0.6740 | 0.0475 | 0.6361 | 0.0850 | 0.5163 | 0.0525 | 0.4110 |
| Binoculars | 0.9900 | 0.9999 | 0.9750 | 0.9952 | 0.9650 | 0.9957 | 0.9125 | 0.9930 |

### v2

@@ 2 Related Work and Background
Evaluated Detectors and Datasets. In our paper, we evaluate seven representative detectors: RADAR (Hu et al., 2023), Detection in the Wild (Wild) (Li et al., 2024), T5Sentinel (Chen et al., 2023), Fast-DetectGPT (Bao et al., 2024), GPTID (Tulchinskii et al., 2023), LogRank (Mitchell et al., 2023), and Binoculars (Hans et al., 2024). RADAR, Wild, and T5Sentinel are trained detectors, while Fast-DetectGPT, GPTID, LogRank, and Binoculars are zero-shot detectors. To ensure a fair comparison and assess the detectors’ ability to generalize to new data, we carefully select datasets that have not been used in the training or evaluation of these detectors. Table 1 presents an overview of the datasets and domains on which each detector has been evaluated. Several datasets, such as Xsum, SQuAD, and Reddit Writing Prompts, have been used in the evaluation or training of multiple detectors. Although these detectors achieve strong Area Under the Receiver Operating Characteristic (AUROC) scores on these datasets, they do not report the True Positive Rate at a set False Positive Rate (TPR@FPR), which is a crucial metric in real-world scenarios. To address this gap, we aim to evaluate all seven detectors on the same datasets using both AUROC and TPR at FPR metrics.
@@ 6 Conclusion
This study evaluates seven advanced detectors across seven tasks and four languages, revealing notable inconsistencies in their detection capabilities. We also examined three different prompting strategies and their impact on detectability, finding that requests for more "human-like" output do not make the text harder to detect, while rewritten human content proves more difficult to identify.
@@ A.1 Results by Model
| GPT-4o | Binoculars | 0.14 | 0.8739 |
| Fast-DetectGPT | 0.07 | 0.8223 | |
| LogRank | 0.00 | 0.5722 | |
| PHD | 0.00 | 0.3976 | |
| Radar | 0.00 | 0.6594 | |
| T5Sentinel | 0.01 | 0.4604 | |
| Wild | 0.00 | 0.4724 | |
| Llama-3 | Binoculars | 0.45 | 0.9474 |
| Fast-DetectGPT | 0.16 | 0.8211 | |
| LogRank | 0.00 | 0.6060 | |
| PHD | 0.04 | 0.4940 | |
| Radar | 0.05 | 0.8088 | |
| T5Sentinel | 0.09 | 0.5358 | |
| Wild | 0.06 | 0.6231 | |
| Mistral | Binoculars | 0.39 | 0.9681 |
| Fast-DetectGPT | 0.16 | 0.9029 | |
| LogRank | 0.06 | 0.4552 | |
| PHD | 0.06 | 0.3104 | |
| Radar | 0.00 | 0.5787 | |
| T5Sentinel | 0.02 | 0.3315 | |
| Wild | 0.07 | 0.5412 | |
| Phi-3 | Binoculars | 0.35 | 0.7825 |
| Fast-DetectGPT | 0.21 | 0.8076 | |
| LogRank | 0.08 | 0.5470 | |
| PHD | 0.14 | 0.4657 | |
| Radar | 0.06 | 0.7700 | |
| T5Sentinel | 0.03 | 0.5209 | |
| Wild | 0.12 | 0.5779 |
@@
| GPT-4o | Binoculars | 0.47 | 0.8676 |
| Fast-DetectGPT | 0.35 | 0.8918 | |
| LogRank | 0.01 | 0.5972 | |
| PHD | 0.01 | 0.4758 | |
| Radar | 0.02 | 0.3474 | |
| T5Sentinel | 0.01 | 0.4909 | |
| Wild | 0.02 | 0.5088 | |
| Llama-3 | Binoculars | 0.66 | 0.9501 |
| Fast-DetectGPT | 0.68 | 0.9465 | |
| LogRank | 0.03 | 0.7816 | |
| PHD | 0.02 | 0.6637 | |
| Radar | 0.13 | 0.6581 | |
| T5Sentinel | 0.03 | 0.4995 | |
| Wild | 0.06 | 0.6024 | |
| Mistral | Binoculars | 0.56 | 0.8954 |
| Fast-DetectGPT | 0.57 | 0.8852 | |
| LogRank | 0.04 | 0.7052 | |
| PHD | 0.04 | 0.6261 | |
| Radar | 0.04 | 0.5776 | |
| T5Sentinel | 0.02 | 0.5078 | |
| Wild | 0.04 | 0.5664 | |
| Phi-3 | Binoculars | 0.48 | 0.8188 |
| Fast-DetectGPT | 0.41 | 0.8424 | |
| LogRank | 0.13 | 0.7448 | |
| PHD | 0.13 | 0.6553 | |
| Radar | 0.07 | 0.5966 | |
| T5Sentinel | 0.02 | 0.5787 | |
| Wild | 0.07 | 0.5923 |
@@
| GPT-4o | Binoculars | 0.05 | 0.6533 |
| Fast-DetectGPT | 0.11 | 0.6981 | |
| LogRank | 0.23 | 0.7070 | |
| PHD | 0.00 | 0.4922 | |
| Radar | 0.00 | 0.2085 | |
| T5Sentinel | 0.01 | 0.4339 | |
| Wild | 0.11 | 0.4753 | |
| Llama-3 | Binoculars | 0.33 | 0.8134 |
| Fast-DetectGPT | 0.19 | 0.7540 | |
| LogRank | 0.61 | 0.8941 | |
| PHD | 0.30 | 0.6642 | |
| Radar | 0.08 | 0.6605 | |
| T5Sentinel | 0.05 | 0.4720 | |
| Wild | 0.15 | 0.6570 | |
| Mistral | Binoculars | 0.06 | 0.6011 |
| Fast-DetectGPT | 0.08 | 0.5862 | |
| LogRank | 0.30 | 0.7635 | |
| PHD | 0.01 | 0.5216 | |
| Radar | 0.00 | 0.3428 | |
| T5Sentinel | 0.02 | 0.4256 | |
| Wild | 0.13 | 0.5931 | |
| Phi-3 | Binoculars | 0.48 | 0.7501 |
| Fast-DetectGPT | 0.30 | 0.7135 | |
| LogRank | 0.60 | 0.9049 | |
| PHD | 0.33 | 0.7668 | |
| Radar | 0.09 | 0.7492 | |
| T5Sentinel | 0.01 | 0.3561 | |
| Wild | 0.38 | 0.8890 |
@@
| GPT-4o | Binoculars | 0.68 | 0.9362 |
| Fast-DetectGPT | 0.52 | 0.9277 | |
| LogRank | 0.37 | 0.7857 | |
| PHD | 0.07 | 0.6279 | |
| Radar | 0.03 | 0.6143 | |
| T5Sentinel | 0.06 | 0.5045 | |
| Wild | 0.01 | 0.5633 | |
| Llama-3 | Binoculars | 0.80 | 0.9695 |
| Fast-DetectGPT | 0.67 | 0.9487 | |
| LogRank | 0.42 | 0.8235 | |
| PHD | 0.24 | 0.6772 | |
| Radar | 0.14 | 0.6907 | |
| T5Sentinel | 0.04 | 0.5265 | |
| Wild | 0.03 | 0.6378 | |
| Mistral | Binoculars | 0.64 | 0.9150 |
| Fast-DetectGPT | 0.56 | 0.9246 | |
| LogRank | 0.33 | 0.7827 | |
| PHD | 0.17 | 0.6514 | |
| Radar | 0.19 | 0.6989 | |
| T5Sentinel | 0.02 | 0.4949 | |
| Wild | 0.02 | 0.6044 | |
| Phi-3 | Binoculars | 0.69 | 0.9243 |
| Fast-DetectGPT | 0.59 | 0.8506 | |
| LogRank | 0.60 | 0.9206 | |
| PHD | 0.37 | 0.7745 | |
| Radar | 0.04 | 0.7626 | |
| T5Sentinel | 0.11 | 0.6134 | |
| Wild | 0.10 | 0.6505 |
@@
| GPT-4o | Binoculars | 0.39 | 0.8584 |
| Fast-DetectGPT | 0.33 | 0.8895 | |
| LogRank | 0.00 | 0.6474 | |
| PHD | 0.00 | 0.2399 | |
| Radar | 0.01 | 0.1890 | |
| T5Sentinel | 0.00 | 0.3086 | |
| Wild | 0.06 | 0.6178 | |
| Llama-3 | Binoculars | 0.71 | 0.9248 |
| Fast-DetectGPT | 0.70 | 0.9352 | |
| LogRank | 0.12 | 0.8505 | |
| PHD | 0.09 | 0.5909 | |
| Radar | 0.29 | 0.6430 | |
| T5Sentinel | 0.16 | 0.6843 | |
| Wild | 0.42 | 0.8185 | |
| Mistral | Binoculars | 0.44 | 0.8672 |
| Fast-DetectGPT | 0.38 | 0.8937 | |
| LogRank | 0.01 | 0.6964 | |
| PHD | 0.01 | 0.3262 | |
| Radar | 0.03 | 0.1770 | |
| T5Sentinel | 0.05 | 0.4642 | |
| Wild | 0.04 | 0.5601 | |
| Phi-3 | Binoculars | 0.38 | 0.5372 |
| Fast-DetectGPT | 0.43 | 0.7147 | |
| LogRank | 0.30 | 0.8252 | |
| PHD | 0.48 | 0.7344 | |
| Radar | 0.42 | 0.8498 | |
| T5Sentinel | 0.01 | 0.3748 | |
| Wild | 0.49 | 0.9189 |
@@
| GPT-4o | Binoculars | 0.69 | 0.9247 |
| Fast-DetectGPT | 0.65 | 0.8847 | |
| LogRank | 0.31 | 0.8778 | |
| PHD | 0.00 | 0.7663 | |
| Radar | 0.12 | 0.8791 | |
| T5Sentinel | 0.02 | 0.5718 | |
| Wild | 0.00 | 0.9249 | |
| Llama-3 | Binoculars | 0.80 | 0.9662 |
| Fast-DetectGPT | 0.72 | 0.9305 | |
| LogRank | 0.65 | 0.9240 | |
| PHD | 0.47 | 0.8589 | |
| Radar | 0.39 | 0.8883 | |
| T5Sentinel | 0.11 | 0.5814 | |
| Wild | 0.30 | 0.9541 | |
| Mistral | Binoculars | 0.79 | 0.9659 |
| Fast-DetectGPT | 0.74 | 0.9358 | |
| LogRank | 0.60 | 0.9249 | |
| PHD | 0.06 | 0.8347 | |
| Radar | 0.37 | 0.9015 | |
| T5Sentinel | 0.07 | 0.6376 | |
| Wild | 0.42 | 0.9574 | |
| Phi-3 | Binoculars | 0.70 | 0.8889 |
| Fast-DetectGPT | 0.49 | 0.8178 | |
| LogRank | 0.40 | 0.8272 | |
| PHD | 0.17 | 0.8137 | |
| Radar | 0.47 | 0.9024 | |
| T5Sentinel | 0.03 | 0.4166 | |
| Wild | 0.25 | 0.9410 |
@@
| GPT-4o | Binoculars | 0.12 | 0.7020 |
| Fast-DetectGPT | 0.05 | 0.6539 | |
| LogRank | 0.02 | 0.6059 | |
| PHD | 0.02 | 0.5611 | |
| Radar | 0.01 | 0.6624 | |
| T5Sentinel | 0.01 | 0.3732 | |
| Wild | 0.13 | 0.5788 | |
| Llama-3 | Binoculars | 0.54 | 0.8739 |
| Fast-DetectGPT | 0.40 | 0.8146 | |
| LogRank | 0.23 | 0.7281 | |
| PHD | 0.08 | 0.6874 | |
| Radar | 0.18 | 0.8910 | |
| T5Sentinel | 0.06 | 0.4902 | |
| Wild | 0.37 | 0.7462 | |
| Mistral | Binoculars | 0.31 | 0.7781 |
| Fast-DetectGPT | 0.14 | 0.7139 | |
| LogRank | 0.05 | 0.6167 | |
| PHD | 0.03 | 0.6115 | |
| Radar | 0.08 | 0.8602 | |
| T5Sentinel | 0.02 | 0.3869 | |
| Wild | 0.19 | 0.6562 | |
| Phi-3 | Binoculars | 0.36 | 0.7671 |
| Fast-DetectGPT | 0.20 | 0.6446 | |
| LogRank | 0.09 | 0.6377 | |
| PHD | 0.30 | 0.7130 | |
| Radar | 0.36 | 0.9726 | |
| T5Sentinel | 0.01 | 0.4116 | |
| Wild | 0.45 | 0.8199 |
@@ A.3 Results by Detector
| Code | Reviews | Abstract | Translation ES | Translation FR | Translation ZH | |||||||
@@
| Radar | 0.0258 | 0.7042 | 0.3358 | 0.8928 | 0.1858 | 0.4647 | 0.1806 | 0.8528 | 0.1475 | 0.8565 | 0.3625 | 0.8307 |
| Fast-DetectGPT | 0.1508 | 0.8385 | 0.6475 | 0.8922 | 0.4608 | 0.8583 | 0.1782 | 0.6337 | 0.0817 | 0.6078 | 0.3333 | 0.7953 |
| Wild | 0.0608 | 0.5537 | 0.2425 | 0.9443 | 0.2517 | 0.7288 | 0.1759 | 0.7433 | 0.0875 | 0.5455 | 0.5358 | 0.9049 |
| PHD | 0.0625 | 0.4169 | 0.1758 | 0.8184 | 0.1467 | 0.4729 | 0.0556 | 0.5264 | 0.0075 | 0.6328 | 0.2283 | 0.7915 |
| LogRank | 0.0367 | 0.5451 | 0.4883 | 0.8884 | 0.1075 | 0.7549 | 0.0509 | 0.4446 | 0.0600 | 0.5900 | 0.1792 | 0.7820 |
| T5Sentinel | 0.0400 | 0.4621 | 0.0575 | 0.5519 | 0.0575 | 0.4580 | 0.0231 | 0.2582 | 0.0025 | 0.3473 | 0.0767 | 0.5286 |
| Binoculars | 0.3317 | 0.8930 | 0.7450 | 0.9364 | 0.4783 | 0.7969 | 0.1944 | 0.6989 | 0.1492 | 0.6754 | 0.5233 | 0.8807 |
@@
| Radar | 0.1225 | 0.7542 | 0.0100 | 0.4832 | 0.0525 | 0.5008 | 0.0400 | 0.3730 |
| Fast-DetectGPT | 0.5450 | 0.9063 | 0.4233 | 0.8522 | 0.6483 | 0.9390 | 0.5808 | 0.8864 |
| Wild | 0.1375 | 0.7437 | 0.0308 | 0.5019 | 0.0183 | 0.4660 | 0.0275 | 0.6180 |
| PHD | 0.0600 | 0.4694 | 0.0792 | 0.6228 | 0.0158 | 0.6371 | 0.0733 | 0.7138 |
| LogRank | 0.0725 | 0.7517 | 0.0725 | 0.6427 | 0.0200 | 0.7741 | 0.1308 | 0.8323 |
| T5Sentinel | 0.0558 | 0.6007 | 0.0075 | 0.4353 | 0.0025 | 0.4131 | 0.0175 | 0.6688 |
| Binoculars | 0.6950 | 0.9271 | 0.5292 | 0.8295 | 0.7250 | 0.9326 | 0.4908 | 0.8785 |
@@
| Radar | 0.0825 | 0.5721 | 0.0125 | 0.3839 | 0.0208 | 0.3966 | 0.2433 | 0.6940 |
| Fast-DetectGPT | 0.2175 | 0.8185 | 0.1408 | 0.6446 | 0.1058 | 0.6169 | 0.1950 | 0.6729 |
| Wild | 0.1250 | 0.7436 | 0.2683 | 0.6027 | 0.2350 | 0.5578 | 0.4392 | 0.8499 |
| PHD | 0.1292 | 0.5907 | 0.1417 | 0.6107 | 0.1617 | 0.5875 | 0.1933 | 0.6631 |
| LogRank | 0.7517 | 0.9705 | 0.4425 | 0.8902 | 0.4217 | 0.8754 | 0.4533 | 0.8317 |
| T5Sentinel | 0.1333 | 0.6275 | 0.0042 | 0.2728 | 0.0208 | 0.3219 | 0.0183 | 0.3828 |
| Binoculars | 0.3792 | 0.7916 | 0.2225 | 0.6866 | 0.1942 | 0.6482 | 0.1333 | 0.6935 |
@@
| Radar | 0.7175 | 0.9344 | 0.0583 | 0.6827 | 0.0417 | 0.5549 | 0.1317 | 0.6120 |
| Fast-DetectGPT | 0.7475 | 0.9508 | 0.4300 | 0.8847 | 0.6158 | 0.9098 | 0.4217 | 0.9011 |
| Wild | 0.0292 | 0.8720 | 0.0775 | 0.6266 | 0.1000 | 0.6035 | 0.0617 | 0.3290 |
| PHD | 0.0425 | 0.6224 | 0.1892 | 0.7168 | 0.1842 | 0.7401 | 0.3835 | 0.7962 |
| LogRank | 0.8583 | 0.9889 | 0.4250 | 0.8824 | 0.4283 | 0.8987 | 0.5208 | 0.8870 |
| T5Sentinel | 0.1708 | 0.6628 | 0.0467 | 0.6010 | 0.0508 | 0.4639 | 0.0217 | 0.3446 |
| Binoculars | 0.7767 | 0.9489 | 0.7700 | 0.9407 | 0.7125 | 0.9393 | 0.6467 | 0.9475 |